OpenStudy (anonymous):

Graph the following inequality: x^2+y^2>49

OpenStudy (anonymous):

do you know what \[x^2+y^2=49\] looks like?

OpenStudy (anonymous):

if the answer is "no" that is fine, i will explain

OpenStudy (anonymous):

Could you please explain? I am stuck. Thanks!

OpenStudy (anonymous):

ok \(x^2+y^2=r^2\) is a circle with center at \((0,0)\) and radius \(r\) in your case you have a circle with center \((0,0)\) and radius \(7\)

OpenStudy (anonymous):



OpenStudy (anonymous):

of course you don't have \(x^2+y^2=49\) you have \(x^2+y^2>49\) so shade the region outside of the circle

OpenStudy (anonymous):

Thank you!